# Notions of entropy for unitary representations

### v2

@@ 3.3 Tracial functionals and vectors
The study of these is motivated by two fundamental sources of examples. First, if $\pi$ is a $d$-dimensional representation with $d$ finite, then the normalized trace on $\mathfrak{B}(\mathbf{C}^{d})$ pulls back to the tracial state $\overline{\mathrm{tr}}_{d}\circ\pi$ on $\mathfrak{A}$. Second, for group C*-algebras, tracial states correspond to characters of the group (beware that some references, including [Dix77], use these terms slightly differently).
@@
is actually tracial on the whole of $\mathfrak{N}$. Clearly $\xi$ is still cyclic for $\mathfrak{N}$. It follows that $\xi$ is also cyclic for $\mathfrak{N}^{\prime}$ [Dix81, Corollary I.6.1], hence also separating for both $\mathfrak{N}$ and $\mathfrak{N}^{\prime}$ [Dix81, Proposition I.1.5], and that $\mathfrak{N}$ and $\mathfrak{N}^{\prime}$ are both finite von Neumann algebra [Dix81, Proposition I.6.9(ii)].
Crucially, if $\xi$ is tracial for the von Neumann algebra $\mathfrak{N}$, then the formula
@@
gives a well-defined involution on the subspace $\mathfrak{N}\xi$. The map $J$ is anti-linear, and it converts inner products to their conjugates as a consequence of the trace identity. In particular, $J$ is an isometry for the restriction of the norm of $H$. It therefore extends by continuity to an involution of the whole of $H$ that has the same properties, which we still denote by $J$. It is called the canonical involution associated to $\mathfrak{N}$ and $\xi$.
These constructions go back to Murray and von Neumann. The theory is recounted with more complete references in [Dix81, Section I.5.1–2] in the alternative framework of Hilbert algebras, which are shown to be equivalent to algebras with tracial cyclic vectors in [Dix81, Section I.6.2]. The following facts can all be found in those sections.
@@
    the map from (iii) preserves $\widetilde{\tau}$ and exchanges the von Neumann subalgebras $\mathfrak{N}$ and $\mathfrak{N}^{\prime}$ of $\mathfrak{B}(H)$. ∎
Parts (ii) and (iv) of this lemma also extend to unbounded operators affiliated to $\mathfrak{N}$ or $\mathfrak{N}^{\prime}$. Indeed, this holds for self-adjoint affiliated operators by taking limits of increasing sequences of bounded operators, and then for general affiliated operators by polar decomposition.
Here is the Radon–Nikodym theorem comparing a positive functional to a trace.
@@
In the setting above, let $\varphi$ be a $\lambda$-normal positive functional. Then there is a unique non-negative self-adjoint operator $T$ affiliated to $\mathfrak{N}^{\prime}$ such that $\xi\in\mathrm{dom}\,T$ and $\varphi$ is associated to $\lambda$ by $T\xi$. ∎
As far as I know, a version of Proposition 3.5 appears first as a special case of [Dye52, Theorem 4]; see also [Seg53b, Theorem 14] and the discussion that surrounds it, or [Dix81, Exercise III.1.14]. The proposition really has two parts, and some references keep them separate. The first part shows that if $\varphi$ is associated to $\lambda^{\oplus\infty}$ then it is actually associated to $\lambda$ itself [Dix81, Theorem III.1.4]. This part already depends on the tracial property of $\xi$. The second part shows that the associating vector in $H$ can be chosen to have the desired form $T\xi$ for a unique $T$ as above; see [Dix81, Exercise III.1.13]. See also [BO08, Appendix F] for an account that serves some more recent applications. A generalization of Proposition 3.5 assuming that $\xi$ is cyclic and separating, but not necessarily tracial, is one of the early steps towards the Tomita–Takesaki modular theory. See, for instance, [Tak03, Chapters VI–VIII] or [KR97, Section 9.2].
@@
$\int_{[0,\infty)}t^{2}\ \langle E(dt)\xi,\xi\rangle=\int_{[0,\infty)}t^{2}\ \widetilde{\tau}E(dt)<\infty.$ (3.7)
This says that $T$ is ‘square-integrable’ as defined in [Seg53b, Subsection 3.4].
We sometimes need to reproduce the action of $T$ using $\mathfrak{N}$ rather than $\mathfrak{N}^{\prime}$. To do this, set $S:=JTJ$, where $J$ is the canonical involution. By Lemma 3.4, this satisfies $\xi\in\mathrm{dom}\,S$ and
@@ 3.4 Fuglede–Kadison determinants
This definition is sometimes extended to other invertible elements $a$ by applying it to $|a|$, but we do not need this. See [FK51, FK52] or [Dix81, Section I.6.11].
If $\mathfrak{N}$ is a von Neumann algebra, $\widetilde{\tau}$ is a normal tracial positive functional on it, and $A\in\mathfrak{N}$ is positive and invertible, then $\Delta A$ can be expressed in terms of the spectral resolution of $A$. Using this expression, we can extend the domains of both $\widetilde{\tau}$ and the resulting determinant $\Delta$ to include suitable classes of affiliated operator. If $T$ is a non-negative operator affiliated to $\mathfrak{N}$, and $E$ is its spectral resolution on $[0,\infty)$, then we call $T$ integrable with respect to $\widetilde{\tau}$ if
@@
(notice that this controls the singularity of $\log$ at $\infty$ but not at 00). If this holds, then we define its Fuglede–Kadison determinant with respect to $\widetilde{\tau}$ to be
@@
The integral here is well-defined by log-integrability, but may take the value $\infty$- ∞, in which case $\Delta T:=0$. The properties of square-integrability (from (3.7)), integrability, and log-integrability are progressively weaker. This extension of the Fuglede–Kadison determinant is covered carefully in [HS07, Section 2].
If $\widetilde{\tau}$ is a state, then $\widetilde{\tau}E$ is a probability measure, and we can apply Jensen’s inequality for expectations of concave functions [Rud87, Theorem 3.3] to the integral of the logarithm in (3.10). This yields an infinitary analog of the det-trace inequality (2.1) when $\widetilde{\tau}$ is a state: for any non-negative integrable operator $T$ affiliated to $\mathfrak{N}$, we have
@@
The expression $\langle E(\cdot)x,x\rangle$ is a Borel measure on $\mathbf{R}$, and the function $t^{2}$ is integrable with respect to it because of the assumption that $x\in\mathrm{dom}\,T$. Therefore the dominated convergence theorem gives $\|Tx-T_{\delta}x\|\to 0$ as $\delta\downarrow 0$.
@@
We now connect these preliminary results with Proposition 3.5. Let $\lambda$ be a representation of $\mathfrak{A}$ with a cyclic tracial vector $\xi$, let $\tau$ be the resulting tracial functional on $\mathfrak{A}$, and let $\widetilde{\tau}$ be its extension to $\lambda(\mathfrak{A})^{\prime\prime}$ as in formula (3.5). Let $\widetilde{\tau}$ also denote the tracial positive functional defined on $\mathfrak{N}:=\lambda(\mathfrak{A})^{\prime}$ by the same formula, and write $\Delta$ for the associated Fuglede–Kadison determinant on log-integrable operators affiliated to either von Neumann algebra. Lastly, let us abbreviate $\Delta(\lambda(a))$ to $\Delta a$ for $a\in\mathfrak{A}$ when the role of $\lambda$ is clear. This function on $\mathfrak{A}$ depends only on $a$.
Now let $\varphi$ be another positive functional on $\mathfrak{A}$ that is $\lambda$-normal, so it can be represented using the operator $T$ affiliated to $\mathfrak{N}$ from Proposition 3.5. The operator $T$ is an equivariant analog of the representation of a positive semidefinite matrix $Q$ as the square of another such matrix $V$. In that situation we have $\det Q=(\det V)^{2}$, and this motivates the following definition.
@@
Then the Fuglede–Kadison determinant of $\varphi$ with respect to $\tau$ is
$\Delta_{\tau}\varphi:=(\Delta T)^{2},$
where $T$ is the operator from Proposition 3.5.
@@
Let $\mathfrak{S}_{\delta}$ be the set of self-adjoint elements $R$ of $\mathfrak{N}$ that satisfy $\delta\leq R\leq\delta^{-1}$. The function $R\mapsto\log R$ is strong-operator continuous on the set $\mathfrak{S}_{\delta}$: for instance, this holds by applying [Con00, Lemma 44.2] to a Lipschitz function that agrees with $\log$ on $[\delta,\delta^{-1}]$. Since $\tau$ is positive and normal, it is ultraweakly continuous [Con00, Theorem 46.4], and this in turn implies strong-operator continuity when restricted to any bounded subset of $\mathfrak{N}$ such as $\mathfrak{S}_{\delta}$ (this follows, for instance, from the explicit description of such functionals in [Con00, Theorem 46.4(d)]). By the continuity of composition, it follows that the expression
$\Delta R=\exp(\tau(\log R))$
is also strong-operator continuous on $\mathfrak{S}_{\delta}$. Now the existence of $a\in\mathfrak{A}$ satisfying $\pi(a)\in\mathfrak{S}_{\delta}$ and also the approximations (3.14) follows from the Kaplansky density theorem as in [Con00, Theorem 44.1(c)]. ∎
As far as I know, the first definition of a Fuglede–Kadison determinant for positive functionals is Arveson’s in [Arv67, Definition 4.3.7]. His setting is functionals on a von Neumann algebra, but the definition itself works for a general C*-algebra without change. His definition is different from ours: his expression is a variational principle. For us this becomes the following proposition. We give two slight variants for convenient application later.
@@
Note Proposition 3.9 requires that $\tau$ be normalized. If we drop this assumption then some additional power-normalization appears in (3.15).
@@
Firstly, in case $\mathfrak{A}=\mathrm{M}_{k}$, Proposition 3.9 is a classical variational principle relating traces and determinants [HJ13, Exercise 7.8.P4].
@@
This time, we can recognize $\Delta\varphi_{\mathrm{ac}}$ as the ‘reversed’ relative entropy (also called Kullback–Leibler divergence) $S(\mu\mid\varphi_{\mathrm{ac}})$, and (3.16) is a classical variational principle for this quantity: see, for instance, [Sim05a, Lemma 2.3.3]. Alternatively, we can allow $\mathfrak{A}=L^{\infty}(\mu)$ for some standard probability space $(\Omega,\mu)$ and arrive at an analogous formula in case $d\varphi=f\cdot d\mu$ for some $f\in L^{1}(\mu)$.
@@
Let $(\Omega,\mu)$ be a standard probability space, let $k$ be a positive integer, and let $\mathfrak{A}$ be $L^{\infty}(\mu;\mathrm{M}_{k})$ (see, for instance, [Con00, Secton 52] for standard properties of this construction). Define a tracial state on $\mathfrak{A}$ by
@@
$\Delta q=\exp\frac{1}{k}\int\log\det q(\omega)\ d\mu(\omega)\qquad(q\in\mathfrak{A}_{+}=L^{\infty}(\mu;\mathrm{M}_{k+})),$
@@
On the other hand, an element of $q\in L^{1}(\mu;\mathrm{M}_{k+})$ may be used to define a positive functional on $\mathfrak{A}$ by setting
@@
Let $\tau$ be associated to its GNS representation $\lambda$ by the cyclic tracial vector $\xi$, and let $\widetilde{\tau}$ be its normal extension to $\lambda(\mathfrak{A})^{\prime\prime}$. Associate $\varphi_{\mathrm{ac}}$ to $\lambda$ by a vector of the form $T\xi$ as in Proposition 3.5. In addition, let $\varphi_{\mathrm{sing}}$ be associated to its GNS representation $\pi_{\mathrm{sing}}$ be the vector $u$. Overall, $\varphi$ is associated to $\pi:=\pi_{\mathrm{sing}}\oplus\lambda$ by the vector $(u,T\xi)$.
@@
Step 2. Let $S:=JTJ$, so this satisfies (3.8) and (3.13). If $a\in A$ is positive, invertible, and satisfies $\Delta a\geq 1$, then
@@
The second equality holds by the multiplicativity of $\Delta$ (see [HS07, Proposition 2.5]), and the second inequality is the infinitary det-trace inequality (3.11) (this is where we need the assumption that $\tau$ is normalized). This proves the inequality $\leq$” on the second line in (3.15).
Step 3. Let $S_{\delta}:=(S\vee\delta)\wedge\delta^{-1}$ as in Lemma 3.6, and define the operator
@@
According to Proposition 3.9, the right-hand side of (3.15) does not depend on $\varphi_{\mathrm{sing}}$ at all. The proof explains why. In forming the operator $R_{\delta}$, we use the first summand to suppress the $\lambda$-singular part of $\pi$ as much as we like. Then Kaplansky’s density theorem gives a close enough approximation $a$ to $R_{\delta}$ so that (i) we retain this feature but also (ii) we pay as little as we like from its Fuglede–Kadison determinant, which is a function of $\lambda(a)$.
Proposition 3.9 expresses $\Delta\varphi_{\mathrm{ac}}$ as an infimum of weak continuous positive linear functionals of $\varphi$. Therefore the general theory of infinite-dimensional convex analysis gives the following standard consequences: see, for example, [Sim93, Theorem I.6.3], or compare [Sim05a, Theorem 2.3.4] or [Wal82, Theorem 9.12].
@@
Arveson’s result [Arv67, Corollary 4.3.3] is a special case of this corollary.
Since the function $\log$ is monotone and continuous on $(0,\infty)$, properties (i) and (iii) above carry over to the expression $\log\Delta\varphi_{\mathrm{ac}}$ as well. As written above, property (ii) does not hold for this expression, but using also the concavity of $\log$ we can still conclude that this expression is concave:
@@
Definition 3.7 is a special case of a construction from quantum statistical mechanics. Consider two faithful normal positive functionals $\omega_{1}$ and $\omega_{2}$ on a von Neumann algebra. Araki’s ‘quantum relative entropy’ ${S(\omega_{1}\mid\omega_{2})}$ is defined by integrating $\log$ against the spectral decomposition of the relative modular operator from $\omega_{1}$ to $\omega_{2}$. His papers [Ara76, Ara76, Ara78] study how it generalizes the quantum relative entropy for states on finite-dimensional matrix algebras. Up to a logarithm and a normalization, this includes Definition 4.9 in case one of those functionals is tracial. This generalizes the fact that the negative log-determinant of a finite-dimensional positive semi-definite matrix $Q$ is equal to the finite-dimensional quantum relative entropy $S(I_{k}\mid Q)$: see, for instance, [Bha07, equation (4.19)]. Similarly, the integral from the right-hand side of Szegő’s limit theorem (1.1) is sometimes written as a classical relative entropy (also called Kullback–Leibler divergence) between two finite measures.

### v3

@@ 3.3 Tracial functionals and vectors
The study of these is motivated by two fundamental sources of examples. First, if $\pi$ is a $d$-dimensional representation with $d$ finite, then the normalized trace on $\mathfrak{B}(\mathbf{C}^{d})$ pulls back to the tracial state $\overline{\mathrm{tr}}_{d}\circ\pi$ on $\mathfrak{A}$. Second, for group C*-algebras, tracial states correspond to characters of the group (beware that some references, including [48], use these terms slightly differently).
@@
is actually tracial on the whole of $\mathfrak{N}$. The right-hand side of (3.5) actually defines a normal positive functional on the whole of $\mathfrak{B}(H)$, and we denote this by $\widetilde{\tau}$ as well. Clearly $\xi$ is still cyclic for $\mathfrak{N}$. It follows that $\xi$ is also cyclic for $\mathfrak{N}^{\prime}$ [49, Corollary I.6.1], hence also separating for both $\mathfrak{N}$ and $\mathfrak{N}^{\prime}$ [49, Proposition I.1.5], and that $\mathfrak{N}$ and $\mathfrak{N}^{\prime}$ are both finite von Neumann algebra [49, Proposition I.6.9(ii)].
Next, if $\xi$ is cyclic and tracial for the von Neumann algebra $\mathfrak{N}$, then the formula
@@
gives a well-defined involution on the subspace $\mathfrak{N}\xi$. The map $J$ is anti-linear, and it converts inner products to their conjugates as a consequence of the trace identity. In particular, $J$ is an isometry for the restriction of the norm of $H$ to $\mathfrak{N}\xi$. It therefore extends by continuity to an involution of the whole of $H$ that has the same properties. We still denote this by $J$, and call it the canonical involution associated to $\mathfrak{N}$ and $\xi$.
These constructions go back to Murray and von Neumann. The theory is recounted with more complete references in [49, Section I.5.1–2] in the alternative framework of Hilbert algebras, which are shown to be equivalent to algebras with tracial cyclic vectors in [49, Section I.6.2]. The following facts can all be found in those sections.
@@
    the map from (iii) preserves the functional $\widetilde{\tau}$ from (3.5) and exchanges the von Neumann subalgebras $\mathfrak{N}$ and $\mathfrak{N}^{\prime}$ of $\mathfrak{B}(H)$. ∎
Here is the Radon–Nikodym theorem that compares a positive functional to a trace:
@@
In the setting above, let $\varphi$ be a $\lambda$-normal positive functional. Then there is a unique non-negative operator $T$ affiliated to $\mathfrak{N}^{\prime}$ such that $\xi\in\mathrm{dom}\,T$ and $\varphi$ is associated to $\lambda$ by $T\xi$. ∎
As far as I know, a version of Proposition 3.5 first appeared as a special case of [52, Theorem 4]; see also [132, Theorem 14] and the discussion that surrounds it, or [49, Exercise III.1.14]. Proposition 3.5 really has two substantial parts, followed by a final simplification. Some references present these parts separately, so let us quickly sketch how they are put together:
  • The first part shows that, if $\varphi$ is $\lambda$-normal, then it is actually associated to $\lambda$ itself as opposed to $\lambda^{\oplus\infty}$. This holds because $\xi$ is separating, as discussed above: see [49, Theorem III.1.4]. This consequence of a separating vector is derived from the fact that, if two von Neumann algebras both have cyclic and separating vectors, then any normal isomorphism between them is spatial [49, Theorem III.1.3].
  • The second part shows that any vector in $H$ has the form $T\xi$ for a unique operator $T$ affiliated to $\mathfrak{N}^{\prime}$; see [49, Exercise III.1.13].
  • To finish the proof, one replaces $T$ with $|T|$. This does not change the associated functional $\varphi$ because the unitary part of the polar decomposition of $T$ commutes with $\lambda$. After this replacement, we may assume that $T$ is non-negative, and having done so $\varphi$ determines $T$ uniquely. This is because $\varphi$ determines its own GNS representation and associating vector uniquely up to unitary equivalence, and such a unitary equivalence then collides with the uniqueness of the polar decomposition of $T$.
@@
$\int_{[0,\infty)}t^{2}\ \widetilde{\tau}E(dt)=\int_{[0,\infty)}t^{2}\ \langle E(dt)\xi,\xi\rangle<\infty.$ (3.7)
This says that $T$ is ‘square-integrable’ as defined in [132, Subsection 3.4].
We sometimes need a variant of $T$ affiliated to $\mathfrak{N}$ rather than $\mathfrak{N}^{\prime}$. To obtain this, set $S:=JTJ$, where $J$ is the canonical involution. By Lemma 3.4, this satisfies $\xi\in\mathrm{dom}\,S$ and
@@ 3.4 Fuglede–Kadison determinants
This definition is sometimes extended to other invertible elements $a$ by applying it to $|a|$, but we do not need this. See [63, 64] or [49, Section I.6.11].
If $\mathfrak{N}$ is a von Neumann algebra, $\widetilde{\tau}$ is a normal tracial positive functional on it, and $A\in\mathfrak{N}$ is positive and invertible, then $\Delta A$ can be expressed in terms of the spectral resolution of $A$. Using this expression, we can extend the domains of both $\widetilde{\tau}$ and the resulting determinant $\Delta$ to include suitable classes of affiliated operators. If $T$ is a non-negative operator affiliated to $\mathfrak{N}$, and $E$ is its spectral resolution on $[0,\infty)$, then we call $T$ integrable with respect to $\widetilde{\tau}$ if
@@
(notice that this controls the singularity of $\log$ at $\infty$ but not at 00). If this holds, then we define the Fuglede–Kadison determinant of $T$ with respect to $\widetilde{\tau}$ to be
@@
The integral here is well-defined by log-integrability, but it may take the value $\infty$- ∞, in which case $\Delta T:=0$. The properties of square-integrability (from (3.7)), integrability, and log-integrability are progressively weaker. This extension of the Fuglede–Kadison determinant is covered carefully in [77, Section 2].
If $\widetilde{\tau}$ is a state, then $\widetilde{\tau}E$ is a probability measure, and we can apply Jensen’s inequality for expectations of concave functions [126, Theorem 3.3] to the integral of the logarithm in (3.10). This yields an infinitary analog of the det-trace inequality (2.1) when $\widetilde{\tau}$ is a state: for any non-negative integrable operator $T$ affiliated to $\mathfrak{N}$, we have
@@
The expression $\langle E(\cdot)x,x\rangle$ is a finite Borel measure on $\mathbf{R}$, and the function $t^{2}$ is integrable with respect to it because of the assumption that $x\in\mathrm{dom}\,T$. Therefore the dominated convergence theorem gives $\|Tx-T_{\delta}x\|\to 0$ as $\delta\downarrow 0$.
@@
We now connect these preliminary results with Proposition 3.5. Let $\lambda$ be a representation of $\mathfrak{A}$ with a cyclic tracial vector $\xi$, let $\tau$ be the resulting tracial functional on $\mathfrak{A}$, and let $\widetilde{\tau}$ be its extension to $\lambda(\mathfrak{A})^{\prime\prime}$ as in formula (3.5). Let $\widetilde{\tau}$ also denote the tracial positive functional defined on $\mathfrak{N}:=\lambda(\mathfrak{A})^{\prime}$ by the same formula. Write $\Delta$ for the associated Fuglede–Kadison determinant on log-integrable operators affiliated to either von Neumann algebra. Lastly, abbreviate $\Delta(\lambda(a))$ to $\Delta a$ for $a\in\mathfrak{A}$ when the role of $\lambda$ is clear.
Let $\varphi$ be another positive functional on $\mathfrak{A}$ that is $\lambda$-normal, so it can be represented using the operator $T$ affiliated to $\mathfrak{N}$ from Proposition 3.5. The operator $T$ is an equivariant analog of the representation of a positive semidefinite matrix $Q$ as the square of another such matrix $V$. In that situation we have $\det Q=(\det V)^{2}$, and this motivates the following definition.
@@
The Fuglede–Kadison determinant of $\varphi$ with respect to $\tau$ is
$\Delta_{\tau}\varphi:=(\Delta T)^{2}.$
@@
Let $\mathfrak{S}_{\delta}$ be the set of self-adjoint elements $R$ of $\mathfrak{N}$ that satisfy $\delta\leq R\leq\delta^{-1}$. The function $R\mapsto\log R$ is strong-operator continuous on the set $\mathfrak{S}_{\delta}$: for instance, this holds by applying [38, Lemma 44.2] to a Lipschitz function that agrees with $\log$ on $[\delta,\delta^{-1}]$. Since $\widetilde{\tau}$ is positive and normal, it is ultraweakly continuous, and this in turn implies strong-operator continuity when restricted to any bounded subset of $\mathfrak{N}$ such as $\mathfrak{S}_{\delta}$: this follows, for instance, from the explicit description of such functionals in [38, Theorem 46.4]. By the continuity of composition, it follows that the expression
$\Delta R=\exp(\widetilde{\tau}(\log R))$
is also strong-operator continuous on $\mathfrak{S}_{\delta}$. Now the existence of $a\in\mathfrak{A}$ satisfying $\pi(a)\in\mathfrak{S}_{\delta}$ and also the approximations (3.14) follows from the Kaplansky density theorem as in [38, Theorem 44.1(c)]. ∎
As far as I know, the first definition of a Fuglede–Kadison determinant for positive functionals was Arveson’s in [9, Definition 4.3.7]. His setting is functionals on a von Neumann algebra, but the definition itself works for a general C*-algebra without change. His definition is different from ours: his expression is a variational principle. For us this becomes the following proposition.
@@
Beware that Proposition 3.9 requires $\tau$ to be normalized.
@@
Firstly, in case $\mathfrak{A}=\mathrm{M}_{k}$, Proposition 3.9 is a classical variational principle relating traces and determinants: see [88, Exercise 7.8.P4], for example.
@@
This time, we can recognize $\Delta\varphi_{\mathrm{ac}}$ as the ‘reversed’ relative entropy (also called Kullback–Leibler divergence) $S(\mu\mid\varphi_{\mathrm{ac}})$, and (3.16) is a classical variational principle for this quantity: see, for instance, [136, Lemma 2.3.3]. Alternatively, we can allow $\mathfrak{A}=L^{\infty}(\mu)$ for some standard probability space $(\Omega,\mu)$ and arrive at an analogous formula in case $\varphi\ll\mu$.
@@
Let $(\Omega,\mu)$ be a standard probability space, let $k$ be a positive integer, and let $\mathfrak{A}$ be $L^{\infty}(\mu;\mathrm{M}_{k})$ (see, for instance, [38, Secton 52] for properties of this algebra). Define a tracial state on $\mathfrak{A}$ by
@@
$\Delta b=\exp\frac{1}{k}\int\log\det b(\omega)\ d\mu(\omega)\qquad(b\in\mathfrak{A}_{+}=L^{\infty}(\mu;\mathrm{M}_{k+})),$
@@
On the other hand, any $q\in L^{1}(\mu;\mathrm{M}_{k+})$ may be used to define a positive functional on $\mathfrak{A}$ by setting
@@
Let $\tau$ be associated to its GNS representation $\lambda$ by the cyclic tracial vector $\xi$, and let $\widetilde{\tau}$ be its normal extension to $\lambda(\mathfrak{A})^{\prime\prime}$ as in (3.5). Associate $\varphi_{\mathrm{ac}}$ to $\lambda$ by a vector of the form $T\xi$ as in Proposition 3.5. Let $S:=JTJ$, so this satisfies (3.8) and (3.13). In addition, let $\varphi_{\mathrm{sing}}$ be associated to its GNS representation $\pi_{\mathrm{sing}}$ by the vector $u$. Overall, $\varphi$ is associated to $\pi:=\pi_{\mathrm{sing}}\oplus\lambda$ by the vector $(u,S\xi)$.
@@
Step 1. If $a\in A$ is positive, invertible, and satisfies $\Delta a\geq 1$, then
@@
The second equality holds by the multiplicativity of $\Delta$ (see [77, Proposition 2.5]), and the second inequality is the infinitary det-trace inequality (3.11) (this is where we need the assumption that $\tau$ is normalized). This proves the inequality $\leq$”.
Step 2. Let $S_{\delta}:=(S\vee\delta)\wedge\delta^{-1}$ as in Lemma 3.6, and define the operator
@@
According to Proposition 3.9, the right-hand side of (3.15) does not depend on $\varphi_{\mathrm{sing}}$ at all. The proof explains why. In forming the operator $R_{\delta}$, we use the first summand to suppress the $\lambda$-singular part of $\pi$ as much as we like. Then Kaplansky’s density theorem gives a close enough approximation $a$ to $R_{\delta}$ so that we retain this feature but also pay as little as we like from its Fuglede–Kadison determinant, which is a function of $\lambda(a)$.
Proposition 3.9 expresses $\Delta\varphi_{\mathrm{ac}}$ as an infimum of weak continuous positive linear functionals of $\varphi$. Therefore the general theory of infinite-dimensional convex analysis gives the following standard consequences: see, for example, [135, Theorem I.6.3], or compare [136, Theorem 2.3.4] or [157, Theorem 9.12].
@@
Arveson’s result [9, Corollary 4.3.3] is a special case of this corollary.
Since the function $\log$ is monotone and continuous on $(0,\infty)$, conclusions (a) and (c) from Corollary 3.11 carry over to the expression $\log\Delta\varphi_{\mathrm{ac}}$ as well. As written above, conclusion (b) does not hold for this expression, but using also the concavity of $\log$ we can still conclude that this expression is concave:
@@
Definition 3.7 is a special case of a construction from quantum statistical mechanics. Consider two faithful normal positive functionals $\omega_{1}$ and $\omega_{2}$ on a von Neumann algebra. Araki’s ‘quantum relative entropy’ ${S(\omega_{1}\mid\omega_{2})}$ is defined by integrating $\log$ against the spectral decomposition of the relative modular operator from $\omega_{1}$ to $\omega_{2}$. His papers [5, 6, 7] studied how it generalizes the quantum relative entropy for states on finite-dimensional matrix algebras. Up to a logarithm and a normalization, this includes Definition 4.12 in case one of those functionals is tracial. This generalizes the fact that the negative log-determinant of a finite-dimensional positive semi-definite matrix $Q$ is equal to the finite-dimensional quantum relative entropy $S(I_{k}\mid Q)$: see, for instance, [16, equation (4.19)]. Similarly, the integral from the right-hand side of Szegő’s limit theorem (1.1) is sometimes written as a classical relative entropy (also called Kullback–Leibler divergence) between two finite measures.